# Bayesian changepoint detection via logistic regression and the topological analysis of image series

### v2

@@ 4.1 Simulation
We compare our method—which we deem bclr [^3] Bayesian Changepoint via Logistic Regression.—to the random forest based classification changepoint detection method from the changeforest package (Londschien et al., 2023) (which we deem cf in the sequel); the E-divisive method from the ecp package (James and Matteson, 2015; Matteson and James, 2014); and, the kcp kernel change-point method from Arlot et al. (2019) as implemented in the ruptures package (Truong et al., 2020). To demonstrate the dual benefit of our approach in conjunction with TDA, we gave our method the TDA features from $f_{\text{stat}}$, and the other methods dimensionality-reduced features by first vectorizing the images and then projecting them down onto their first 36 principal components. We did consider a Bayesian changepoint method with available code, but this did not perform well. Information about this and the parameters used for each method can be seen in Section S2 of the Supplementary Material.
We also devised new topologically-aware versions of cf, ecp, and kcp—where we fed these algorithms the same topological features from $f_{\text{stat}}$. We christened these methods cf+TDA, ecp+TDA, and kcp+TDA respectively. For fair comparison, each algorithm received either the exact same PCA or standardized TDA features (depending on which they were designed for). We estimated the changepoint $\hat{\kappa}$ via the posterior mode in our setup, in light of the theory in Section 2.2 and to ensure the estimate corresponded to an actual frame. For experiment 1, we specified our prior to be $N(0,3I_{d})$ for $\bm{\beta}$ and discrete uniform on $\{1,\dots,50\}$ for $\kappa$. For our method, we always took the 2500 posterior samples for both $\kappa$ and $\bm{\beta}$ after the burn-in period (more details on the properties of the Gibbs sampler and its convergence are available in Section S2.2 of the Supplementary Material).
@@
To verify that our bclr method performs well for more than just two specific videos, we generated 1000 additional i.i.d. random image series $V^{(1)}$ (Experiment 1) and $V^{(2)}$ (Experiment 2) and gauged the behavior of the various methods in both experiments. For each of the methods, we calculated the accuracy of the estimated changepoint $\hat{\kappa}$ in terms of proportion of times the estimated changepoint was exactly correct (“Exact”) and the root mean-squared error, or “RMSE” (of the estimated changepoint from $\kappa^{*}=25,40$). For our bclr method we calculated the RMSE within each of the posterior samples of length 2500 and report the mean RMSE across all 1000 simulations. Because we calculate the RMSE in this way, we are able to report the standard error of the RMSE. In Section S2.4 in the Supplementary Material, we expound on this line of thought and also report the RMSE of the posterior mean and the posterior mode for a few select cases.
@@
Even though $\kappa$ is discrete, the credible intervals for the TDA features in our setup are conservative at each setting of $\alpha$ that we consider. Though the specified intervals do not even necessarily have the highest posterior mass (we do this in Table S1 and see similar results), Table 2 indicates that interval estimation using bclr with TDA features is appropriate here (where the signal is fairly strong relative to the noise), in a way that imposes no distributional restrictions on the data $\bm{X}$. A more detailed analysis of these intervals can be seen in Section S3 of the Supplementary Material. For Experiment 2 ($\kappa^{*}=40$) the signal is cut in half, and corresponding the coverage probabilities [^4] The study of frequentist coverage probability bias of these types of Bayesian posterior quantile intervals was studied in Sweeting (2001), albeit those results are not applicable to this more complicated setting. are less than their nominal amount—see Table S2. Nevertheless, the probability our method will return an estimate containing the true changepoint is higher than the other methods do to our ability to seemlessly carry out interval estimation.
@@
so that the unique mode of $q$ is $\kappa=40$ and no prior mass is placed on $\kappa=50$. For this experiment we also conducted 5000 Monte Carlo iterations and chose the 2500 simulations after the burn-in. Additional details on the performance for this method with alternative priors for $\bm{\beta}$ and $\kappa$ can be seen in Section S2.1 in the Supplementary Material. As one can see in Table 3, our method using the above specified priors outperforms both kcp+TDA and ecp+TDA methods and vastly outperforms the non-topological versions of those methods as well as all versions of cf.
@@
Furthermore, none of other methods we describe tell us what coordinates are most important for detecting such a change. With regards to the 1000 simulated videos in Experiment 1 ($\kappa^{*}=25$) we measured the importance of a given coordinate of $\bm{\beta}$ with respect to the TDA features via the mean and signal-to-noise ratio (SNR) [^5] This quantity is meaningful as each coordinate has mean 0 and variance 1. It is defined as $\bar{\beta_{j}}^{2}/s(\beta_{j})^{2}$, $j=1,\dots,d$.. We can see the mean and standard deviation of the SNRs—for persistence statistics which had the highest SNR most often among the 1000 simulations—in Table 4. Beyond this, we also were able to get the mean posterior correlations for the regression coefficients corresponding to the persistence statistics. The largest absolute mean correlation among the 5 statistics in Table 4 was -0.358 between kurtosis and skewness of the lifetimes for $\mathcal{D}_{0}$. The smallest absolute mean correlation was -0.007 between the ALPS statistic and the persistent entropy. The other absolute mean correlations hovered between 0.043 and 0.2, which indicates that there was not necessary a single statistic which stood out above the rest and justified the inclusion of multiple statistics and thus the feature embedding $f_{\text{stat}}$.
The ability of our method to determine which coordinate is important has further utility than the detection of a topological change. In Section 5, we show the efficacy of our method in detecting where and how a change in data of mixed type occurs. Furthermore, we demonstrate that our method can not only detect a change in covariance via a degree-2 polynomial feature embedding as well as any other advanced nonparametric method—but also identify the nature of said change.
@@
persistent entropy of $l_{i}$ for $\mathcal{D}^{0}$ 0.185 1.562 (1.082)
skewness of $l_{i}$ for $\mathcal{D}^{0}$ 0.156 1.873 (0.817)
kurtosis of $l_{i}$ for $\mathcal{D}^{0}$ 0.120 1.734 (0.875)
ALPS statistic of $l_{i}$ for $\mathcal{D}^{0}$ 0.114 1.118 (1.218)
variance of $l_{i}$ for $\mathcal{D}^{0}$ 0.111 1.703 (0.683)
@@
4.2 Applications
Evidence presented in Thomas et al. (2023) indicates that a reasonable topological summary for the detection of nanoparticle dynamics is some linear combination of the ALPS statistic (ibid.) and persistent entropy (Atienza et al., 2020). We assess the accuracy of this statement using the topological embedding $f_{\text{stat}}$. Visual inspection of the nanoparticle video of interest [^6] Described in detail in the Supplementary Material of Thomas et al. (2023). suggests that a change occurs in the vicinity of frame 210, though the high signal-to-noise ratio precludes any notion of “ground truth”. Thus, we apply our algorithm to the data to get an estimate of where the change occurs and what the best representation for said change is. Plots describing these results can be seen in Figure 3 and averages of 5 consecutive frames (to improve visualization) before and after the estimated changepoint can be seen in Figure 4. We chose a prior for $\bm{\beta}$ with mean 0 and covariance matrix equal to $3I_{d}$, where $I_{d}$ is the $d$-dimensional identity matrix.
As seen in Figure 3, the marginal posterior distribution of $\kappa$ concentrates around frame 210. After having run our Gibbs sampler for 5000 iterations and discarding the first 2500 samples from the posterior Monte Carlo sample, we estimate that $\pi(210\mid\bm{X})=0.3560$, $\pi(211\mid\bm{X})=0.3112$, and $\pi(214\mid\bm{X})=0.2984$ with probability less than $0.02$ elsewhere. We can see using the posterior mean of $\bm{\beta}$ that reasonable separation is achieved for $\bm{x}_{\kappa}^{\top}\bm{\beta}$ in Figure 3 (bottom). The question still remains as to which topological statistics best represent the change. We summarize the importance of a given coordinate of $\bm{\beta}$ by its signal-to-noise ratio (SNR), defined in the previous section. The top 5 statistics in terms of signal-to-noise ratio can be seen in Table 5.
Statistic Posterior SNR Posterior mean
$\bm{25^{th}}$ percentile of $l_{i}$ for $\mathcal{D}^{1}$ 5.694 1.087
variance of $l_{i}$ for $\mathcal{D}^{0}$ 3.772 2.389
persistent entropy of $l_{i}$ for $\mathcal{D}^{1}$ 3.407 -1.481
persistent entropy of $l_{i}$ for $\mathcal{D}^{0}$ 2.117 -0.905
ALPS statistic of $l_{i}$ for $\mathcal{D}^{0}$ 1.890 -1.222
The results of Table 5 support—but also refine—the results of Thomas et al. (2023), wherein the persistent entropy and the ALPS statistic of the lifetimes of $\mathcal{D}^{0}$ were chosen to represent the dynamics of the nanoparticle. These results seem to suggest that the variance of $l_{i}$ of $\mathcal{D}^{0}$ captures the dynamics even better, having a large posterior mean and rather low variance. For the sake of comparison, we ran ecp+TDA, cf+TDA, and kcp+TDA and recovered an estimated changepoint of $\hat{\kappa}=214$ for all three methods with $p$-values of 0.005 for the first two.
We conclude this section by looking at a 100 frame solar flare video, which was taken from Xie et al. (2012) and analyzed via the topological online changepoint method PERCEPT proposed in Zheng et al. (2023). Frames that depict the video immediately prior to and after then changepoint can be seen in Figure 6. Running our changepoint method with the same prior as in the case of the nanoparticle video, we observe a posterior mode of $\hat{\kappa}=44$ using bclr with $f_{\text{stat}}$ (see histogram (top) in Figure 5). However, using the pixel means of the images as a one-dimensional signal to detect the solar flare [^7] Here we do not standardize each frame to have mean zero intensity. the estimated posterior mode of $\pi(\kappa\mid\bm{X})$ is $\kappa=46$. This seems reasonable given the below results for the other methods. In the article (Zheng et al., 2023), the changepoint was estimated to be $\kappa=49$ using PERCEPT. Since they were using a CUSUM approach on their derived topological features, their estimate was more likely to occur after the best separation between pre- and post-change distributions. Also, there is no indication in the article that their images were standardized, which leads one to believe that their estimated changepoint captures mostly the increase in the mean intensity of the solar flare video. That being said, using the PCA features described below, our method has posterior mode $\hat{\kappa}=49$. The discrepancy between the estimate of the changepoint using topological statistics and pixel means suggests that there is topological change occurring prior to the solar event which ultimately takes place and leads to a great increase in radiation which is visually apparent.
5 Simulation studies: mixed data and a change in covariance
@@ 5.1 Detecting a change in data of mixed type
To evaluate the performance of bclr in detecting and characterizing a change in data of mixed type, we simulated 2500 independent sequences of the form
$X_{1},\dots,X_{600},$
where $X_{i}=(X_{i1},X_{i2},X_{i3},X_{i4},X_{i5})$, $i=1,\dots,600$ are five-dimensional. The coordinates of $X_{i}$ are defined by
$(X_{i3},X_{i4},X_{i5})=(Y_{i3},Y_{i4},Y_{i5})B$
where each $Y$-coordinate are i.i.d. standard Laplace random variables and $B$ is the matrix defined by
@@
We chose the Laplace distribution to see how our method would fare in the presence of heavier-tailed noise. Coordinates $X_{i1}$ and $X_{i2}$ correspond to a dummy coding of categorical variables $Z_{i}\in\{1,2,3\}$ with distribution $\mathbb{P}_{0}$ until frame 350 and distribution $\mathbb{P}_{1}$ after frame 350—where
@@
Table 6: Comparison of various changepoint detection methods in the discrete variable change setup. Here raw indicates the method was fed data that was NOT centered to have mean zero and standardized to have standard deviation 1 in each coordinate. Standard errors for % Exact and RMSE indicated in parentheses. SEs for RMSE not applicable for the other methods.
Figure 7: Distribution of the posterior means of $\bm{\beta}$ (on standardized scale) over all 2500 simulations for the mixed data change setup of Section 5.1. This figure clearly indicate a strong influence of the component of $\bm{\beta}$ associated to a drop in the value of $P(Z=1)$ after the change (via $x_{1}$).
@@ 5.2 Detecting a change in covariance
To conclude our experiments, we study the rather difficult problem of detecting a change in covariance. As in the previous subsection, we simulated 2500 independent sequences
$X_{1},\dots,X_{300},$
@@
so that $\psi:\mathbb{R}^{4}\to\mathbb{R}^{14}$. Per usual, we standardize [^8] That is, subtract from each column the mean of the series and the divide by the series’ standard deviation. each series $(\psi(X_{i}),i=1,\dots,300)$ prior to feeding it to our algorithm. For our method, we used the same priors as Section 5.1 and ran the algorithm for 5000 iterations, discarding the first 2500 as burn-in. We compare our bclr method to five alternatives and use the same parameters for these methods as the previous section. We apply all three of these methods to the standardized data. We also applied cf and a gaussian likelihood-based method (Lavielle and Teyssiere, 2006, which we deem gaussian) to the raw data $(X_{i},i=1,\dots,300)$ as a basis of comparison. The results of the experiment for all methods are seen in Table 7, and the boxplot of the posterior means of the $\bm{\beta}$ coefficients from our method are seen in Figure 8. We see that our method greatly outperforms all but the gaussian changepoint method, which ought to work well as it satisfies the parametric assumptions of the problem.
Figure 8: Distribution of the posterior means of $\bm{\beta}$ (on standardized scale) over all 2500 simulations for the covariance change setup of Section 5.2. This figure clearly indicate a strong influence of the component of $\bm{\beta}$ associated to the product term $x_{1}x_{2}$, which correctly suggests an increase in the correlation between the first two coordinates of $X_{i}$.

### v3

@@ 4.1 Detecting a topological change in image series
We compare our method—which we deem BCLR [^3] Bayesian Changepoint via Logistic Regression.—to the random forest based classification changepoint detection method from the changeforest package (Londschien et al., 2023) (which we deem CF in the sequel); the E-divisive method (ECP) from the ecp package (James and Matteson, 2015; Matteson and James, 2014); and, the KCP kernel change-point method from Arlot et al. (2019) as implemented in the ruptures package (Truong et al., 2020). To demonstrate the dual benefit of our approach in conjunction with TDA, we gave our method the TDA features from $f_{\text{stat}}$, and the other methods dimensionality-reduced features by first vectorizing the images and then projecting them down onto their first 36 principal components. We did consider a Bayesian changepoint method with available code, but this did not perform well. Information about this and the parameters used for each method can be seen in Section S2 of the Supplementary Material.
We also devised new topologically-aware versions of CF, ECP, and KCP—where we fed these algorithms the same topological features from $f_{\text{stat}}$. We christened these methods CF+TDA, ECP+TDA, and KCP+TDA respectively. For fair comparison, each algorithm received either the exact same PCA or standardized TDA features (depending on which they were designed for). We estimated the changepoint $\hat{\kappa}$ via the posterior mode in our setup, in light of the theory in Section 2.2 and to ensure the estimate corresponded to an actual frame. For experiment 1, we specified our priors to be $N(0,3I_{d})$ for $\bm{\beta}$ and discrete uniform on $\{1,\dots,50\}$ for $\kappa$. For our method, we always took the 2500 posterior samples for both $\kappa$ and $\bm{\beta}$ after the burn-in period of 2500 iterations (more details on the properties of the Gibbs sampler and its convergence are available in Section S2.2 of the Supplementary Material).
@@
To verify that our BCLR method performs well for more than just two specific videos, we generated 1000 additional i.i.d. random image series $V^{(1)}$ (Experiment 1) and $V^{(2)}$ (Experiment 2) and gauged the behavior of the various methods in both experiments. For each of the methods, we calculated the accuracy of the estimated changepoint $\hat{\kappa}$ in terms of proportion of times the estimated changepoint was exactly correct (“Exact”) and the root mean-squared error, or “RMSE” (of the estimated changepoint from $\kappa^{*}=25,40$). For our BCLR method we calculated the RMSE within each of the posterior samples of length 2500 and reported the mean RMSE across all 1000 simulations. Because we calculated the RMSE in this way, we are able to report the standard error of the RMSE. In Section S2.4 in the Supplementary Material, we expound on this line of thought and also report the RMSE of the posterior mean and the posterior mode for a few select cases.
@@
Even though $\kappa$ is discrete, the credible intervals for the TDA features in our setup are conservative at each setting of $\alpha$ that we consider. Though the specified intervals do not necessarily have the highest posterior mass (we do this in Table S1 and see similar results), Table 2 indicates that interval estimation using BCLR with TDA features is appropriate here (where the signal is fairly strong relative to the noise), in a way that imposes no distributional restrictions on the data $\bm{X}$. A more detailed analysis of these intervals can be seen in Section S3 of the Supplementary Material. For Experiment 2 ($\kappa^{*}=40$) the signal is cut in half, and corresponding the coverage probabilities [^4] The study of frequentist coverage probability bias of these types of Bayesian posterior quantile intervals was studied in Sweeting (2001), albeit those results are not applicable to this more complicated setting. are less than their nominal amount—see Table S2. Nevertheless, the probability our method will return an estimate containing the true changepoint is higher than the other methods due to our ability to seamlessly carry out interval estimation.
@@
so that the unique mode of $q$ is $\kappa=40$ and no prior mass is placed on $\kappa=50$. For this experiment, we also conducted 5000 Monte Carlo iterations and chose the 2500 simulations after the burn-in. Additional details on the performance of this method with alternative priors for $\bm{\beta}$ and $\kappa$ can be seen in Section S2.1 of the Supplementary Material. As one can see in Table 3, our method using the priors specified above outperforms both KCP+TDA and ECP+TDA methods and vastly outperforms the non-topological versions of those methods as well as all versions of CF.
@@
Furthermore, none of other methods we describe tell us what coordinates are most important for detecting such a change. With regard to the 1000 simulated videos in Experiment 1 ($\kappa^{*}=25$) we measured the importance of a given coordinate of $\bm{\beta}$, with respect to the TDA features, via the signal-to-noise ratio (SNR) [^5] This quantity is meaningful as each coordinate has mean 0 and variance 1. It is defined as $\bar{\beta_{j}}^{2}/s(\beta_{j})^{2}$, $j=1,\dots,d$.. We can see the mean and standard deviation of the SNRs—for persistence statistics which had the highest SNR most often among the 1000 simulations—in Table 4. We were also able to get the mean posterior correlations for the regression coefficients corresponding to the persistence statistics. The largest absolute mean correlation among the 5 statistics in Table 4 was -0.358 between kurtosis and skewness of the lifetimes for $\mathcal{D}_{0}$. The smallest absolute mean correlation was -0.007 between the ALPS statistic and the persistent entropy. The other absolute mean correlations hovered between 0.043 and 0.2, indicating that there was not necessarily a single statistic that stood out above the rest and justified the use of the feature embedding $f_{\text{stat}}$.
@@
persistent entropy of $l_{p}$ for $\mathcal{D}^{0}$ 0.185 1.562 (1.082)
skewness of $l_{p}$ for $\mathcal{D}^{0}$ 0.156 1.873 (0.817)
kurtosis of $l_{p}$ for $\mathcal{D}^{0}$ 0.120 1.734 (0.875)
ALPS statistic of $l_{p}$ for $\mathcal{D}^{0}$ 0.114 1.118 (1.218)
variance of $l_{p}$ for $\mathcal{D}^{0}$ 0.111 1.703 (0.683)
@@ 4.2 Detecting a change in data of mixed type
To evaluate the performance of BCLR in detecting and characterizing a change in data of mixed type, we simulated 2500 independent sequences of the form $X_{1},\dots,X_{600}$ where each observation $X_{i}=(X_{i1},X_{i2},X_{i3},X_{i4},X_{i5})$, $i=1,\dots,600$ is five-dimensional. The coordinates of $X_{i}$ are defined by $(X_{i3},X_{i4},X_{i5})=(Y_{i3},Y_{i4},Y_{i5})B$ where each of the $Y$-coordinates are i.i.d. standard Laplace random variables and $B$ is the matrix defined by
@@
We chose the Laplace distribution to see how our method would fare in the presence of heavier-tailed noise and we chose $B$ to induce dependence between coordinates 4 and 5, to make the changepoint task more difficult. Coordinates $X_{i1}$ and $X_{i2}$ correspond to a dummy coding of categorical variables $Z_{i}\in\{1,2,3\}$ with distribution $\mathbb{P}_{0}$ until frame 350 and distribution $\mathbb{P}_{1}$ after frame 350—where
@@
Table 5: Comparison of various changepoint detection methods in the discrete variable change setup. Here “raw” indicates the method was fed data that was NOT centered to have mean zero and standardized to have standard deviation 1 in each coordinate. Standard errors for % Exact and RMSE indicated in parentheses. SEs for RMSE not applicable for the other methods.
Figure 3: Distribution of the posterior means of $\bm{\beta}$ (on standardized scale) over all 2500 simulations for the mixed data change setup of Section 4.2. This figure clearly indicates a strong influence of the component of $\bm{\beta}$ associated with a drop in the value of $P(Z=1)$ after the change (via $x_{1}$).
@@ 4.3 Detecting a change in covariance
so that $\psi:\mathbb{R}^{4}\to\mathbb{R}^{14}$. Per usual, we standardize [^6] That is, subtract from each column the mean of the series and then divide by the series’ standard deviation. each series $(\psi(X_{i}),i=1,\dots,300)$ prior to feeding it to our algorithm. For our method, we used the same priors as Section 4.2 and ran the algorithm for 5000 iterations, discarding the first 2500 as burn-in. We compare our BCLR method to five alternatives and use the same parameters for these methods as the previous section. We apply all three of these methods to the standardized data. We also applied CF and a gaussian likelihood-based method (Lavielle and Teyssiere, 2006, which we deem GAUSSIAN) to the raw data $(X_{i},i=1,\dots,300)$ as a basis of comparison. The results of the experiment for all methods are seen in Table 6, and the boxplot of the posterior means of the $\bm{\beta}$ coefficients from our method are seen in Figure 4. We see that our method greatly outperforms all but the GAUSSIAN changepoint method, which ought to work well as it satisfies the parametric assumptions of the problem.
Figure 4: Distribution of the posterior means of $\bm{\beta}$ (on standardized scale) over all 2500 simulations for the covariance change setup of Section 4.3. This figure clearly indicates a strong influence of the component of $\bm{\beta}$ associated with the product term $x_{1}x_{2}$, which correctly suggests an increase in the correlation between the first two coordinates of $X_{i}$.